OpenStudy (anonymous):

A jar contains 5 blue marbles, 8 red marbles, 4 white marbles, and 3 purple marbles. Suppose you pick a marble at random without looking. Find P (red) Write your answer as a fraction in simplest form.

OpenStudy (anonymous):

i feel it is 2/5

OpenStudy (agreene):

# of interest / # total \[\frac{8}{5+8+4+3}=\frac{8}{20}=\frac{2}{5}\]